### v2

@@ 2.3 An impossibility result
Similar to the classic “impossibility results” (Pfanzagl, 1998; Hirano and Porter, 2012), one may conjecture that there cannot exist any statistical rule that approximates the constrained optimum $g^{\ast}_{P}$ uniformly well in the neighbourhood of $P_{0}$, where $g^{\ast}_{P_{h_{n}}}$ can differ arbitrarily from $g^{\ast}_{P_{0}}$ even though the underlying distributions $P_{h_{n}}$ are arbitrarily close to $P_{0}$. This section formalizes this impossibility result that no statistical rule can be both uniformly asymptotically welfare-efficient and uniformly asymptotically feasible in a sufficiently rich class of distributions $\mathcal{P}$. The point of discontinuity $P_{0}$ is characterized in Assumptions 2.2 and 2.3.
@@
Contiguity. There exists a distribution $P_{0}\in\mathcal{P}$ under which a non-empty set of eligibility policies satisfying the constraint exactly $\mathcal{G}_{0}=\left\{g:B(g;P_{0})=k\right\}$. Furthermore, the class of distributions $\mathcal{P}$ includes a sequence of data distributions $\{P_{h_{n}}\}$ contiguous to $P_{0}$, under which for all $g\in\mathcal{G}_{0}$, there exists some $C>0$ such that
@@
Assumption 2.1 assumes there exists a sequence of data distributions $\{P_{h_{n}}\}$ that differ only marginally relative to $P_{0}$. Moreover, the budget function $B(g;P_{h_{n}})$, evaluated at polices that meet the budget constraint exactly under $P_{0}$, approaches $B(g;P_{0})$ from above. In Online Appendix B.2, I give more primitive assumptions under which Assumption 2.1 is guaranteed to hold, requiring the sequence to be differentiable in quadratic mean at $P_{0}$, and along the sequence, the budget function $B(g;P_{h_{n}})$ is twice continuously differentiable at $P_{0}$ with positive derivatives. These primitive assumptions are relatively weak, and have also been considered in the literature to construct the local parametrization around $P_{0}$, e.g. Hirano and Porter (2009) .
@@
The impossibility result is relevant in real-world settings. It is possible to encounter situations where an expansion of a welfare program has no impact on the budget but leads to a strictly positive increase in welfare, as described in Assumptions 2.2 and 2.3. While the nominal cost of welfare program eligibility may be positive, the Marginal Value of Public Funds (MVPF) framework emphasizes that the relevant cost is the net cost, which incorporates fiscal externalities, such as increased tax revenue if the program increases individuals’ incomes. Specifically, Hendren and Sprung-Keyser (2020) estimate fourteen welfare programs (out of 133) to have negative or zero net cost to the government, which implies these programs “pay for themselves”, aligning with Assumptions 2.2 and 2.3. As a result, the impossibility result suggests that statistical rules designed for such cases may be highly sensitive to small changes in the underlying distributions even in large samples.
@@ 3 Sample-analog rule
I first describe the EWM approach (Kitagawa and Tetenov, 2018) and its direct extension. Since $(\Gamma,R)$ involves potential outcomes, they are often unobserved and require estimation based on RCT that introduces estimation errors in addition to sampling errors. Section 3.2 describes how to construct individuals’ benefit and cost estimates $(\Gamma_{i}^{\ast},R_{i}^{\ast})$. To highlight the drawback of the direct extension to EWM, I first consider settings where we observe an experimental data of sample size $n$ where $(\Gamma,R)$ is directly observable, i.e. $(\Gamma_{i}^{\ast},R_{i}^{\ast})=(\Gamma_{i},R_{i})$. The goal of the simplification is to highlight that the non-uniformity I show below can arise from sampling errors alone.
@@
$\widehat{W}_{n}(g)\coloneqq\frac{1}{n}\sum_{i}\Gamma_{i}^{\ast}\cdot g(X_{i}),\ \widehat{B}_{n}(g)\coloneqq\frac{1}{n}\sum_{i}R_{i}^{\ast}\cdot g(X_{i}).$ (8)
@@ 3.1 Welfare inefficiency of the sample-analog rule
A key insight from Kitagawa and Tetenov (2018) is that without a constraint, the sample-analog rule is uniformly asymptotically welfare-efficient. Unfortunately this intuition does not extend to the current setting where the constraint involves an unknown cost. There are common data distributions under which the amount of welfare loss does not vanish even as the sample size gets larger. Consider a one-dimensional policy class $\mathcal{G}=\{g:g(x)=\mathbf{1}\{x\leq t\}\}$, which is based on thresholds of a one-dimensional continuous characteristic $X$. Suppose the policymakers know benefit is positive for everyone so that welfare function is strictly increasing, and only need to estimate whether a given threshold satisfies a capacity constraint due to imperfect take-up. Furthermore, suppose the experiment sample observes take-up $R_{i}$ so that the only uncertainty arises from sampling errors. Proposition 3.1 shows settings with some zero take-ups satisfy Assumptions 2.2 and 2.3, and the sample-analog rule is neither pointwise asymptotically welfare efficient nor pointwise asymptotically feasible.
@@
One-dimensional threshold and imperfect take-up. Consider the special case where under distribution $P$, the benefit $\Gamma>0$ almost surely and the cost $R\in\{0,1\}$ is binary. Suppose $R=0$ for $X\in[\underline{t},\overline{t}]$, but $Pr_{P}\{R=1\mid X\}\in(0,1)$ otherwise. Then $P$ also satisfies Assumptions 2.2 and 2.3. As $n\rightarrow\infty$, the sample-analog rule $\widehat{g}_{\text{sample}}$ would violate the budget constraint half the time, and incur at least $\epsilon>0$ loss in welfare the other half the time:
$Pr_{P^{n}}\left\{W(g_{P}^{*};P)-W(\widehat{g}_{\text{sample}};P)\geq\epsilon\right\}\rightarrow 50\%.$
@@
The more problematic case is illustrated in Figure 3.1 panel (b), where the sample-analog rule picks a suboptimal threshold because the sample-analog constraint is violated at the constrained optimum $g_{P}^{\ast}$. In the setup of Proposition 3.1, when the sample-analog rule $\widehat{g}_{\text{sample}}$ misses $g_{P}^{\ast}$, it is guaranteed to select a suboptimal policy and therefore it is welfare-inefficient under $P$ even asymptotically. This is because the welfare function strictly increases when the budget function remains constant in the neighborhood of the budget constraint. I demonstrate that the same issue extends to broader contexts by employing a simulation calibrated to a real-world DGP from the OHIE in Section 5.
@@ 3.2 Estimates for benefit and cost
In the remainder of the paper, I present constructive results that modify the sample-analog rule and propose a new rule. To lay the groundwork, this section outlines the benefit and cost estimates. The appropriate expressions for these estimates depend on the type of observed data. Below I state the estimates formed based an RCT that randomly assigns the eligibility, which is the leading case of Kitagawa and Tetenov (2018). The observed data $\{A_{i}^{\ast}\}_{i=1}^{n}$ consists of i.i.d. observations $A_{i}^{\ast}=(Y_{i},Z_{i},D_{i},X_{i})\in\mathcal{A}^{\ast}$. The distribution of $A_{i}^{\ast}$ is induced by the distribution of $(Y_{1},Y_{0},R,X)$ as in the population, as well as the sampling design of the RCT. Here $D_{i}$ is an indicator for being in the eligibility arm of the RCT, $Y_{i}$ is the observed outcome and $Z_{i}$ is the observed cost of providing eligibility to an individual participating in the RCT. The observed cost is mechanically zero if an individual is not randomized into the eligibility arm. The estimates for $(\Gamma,R)$ are
$\Gamma_{i}^{\ast}=\alpha(X_{i},D_{i})\cdot Y_{i},\ R_{i}^{\ast}=\frac{D_{i}}{p(X_{i})}\cdot Z_{i},$ (10)
@@
$\left\{\sqrt{n}\cdot\left(\frac{1}{n}\sum_{i}\Gamma_{i}^{\ast}\cdot g(X_{i})-W(g;P)\right)\right\}_{g\in\mathcal{G}}$ $\rightarrow_{d}G_{P}^{W}$
$\left\{\sqrt{n}\cdot\left(\frac{1}{n}\sum_{i}R_{i}^{\ast}\cdot g(X_{i})-B(g;P)\right)\right\}_{g\in\mathcal{G}}$ $\rightarrow_{d}G_{P}^{B}$
@@
Online Appendix B.3 gives primitive assumptions under which Assumption 3.1 is guaranteed to hold for $\widehat{W}_{n}(\cdot)$ and $\widehat{B}_{n}(\cdot)$ constructed using an RCT such as in (10) or an observational study, assuming unconfoundedness and strong overlap. As standard in the literature, I need to restrict the complexity of the policy class $\mathcal{G}$. The policy class of income thresholds considered in this paper is in fact a rather simple class with VC-dimension $d+1$ where $d$ is the number of different thresholds as in (4).
@@ 3.3 Modification that ensures feasibility
Let $\hat{\mathcal{G}}=\{g\in\mathcal{G}:\widehat{B}_{n}(g)\leq k\}$ denote the set of policies that $\widehat{g}_{\text{sample}}$ can choose from, which contains policies that do not violate the sample-analog of the budget constraint. Unsurprisingly, for any finite sample, $\hat{\mathcal{G}}$ can always contain infeasible policies, sometimes of sizable budget violations. The probability that $\hat{\mathcal{G}}$ contains a policy that violates the population budget constraint by a fixed amount $c$ is the follows:
$Pr_{P^{n}}\left\{\exists g:g\in\hat{\mathcal{G}}\ and\ B(g;P)>k+c\right\}$ (11)
@@
A simple modification to the sample-analog rule can reduce the probability of selecting infeasible policies. Instead of $\hat{\mathcal{G}}$, let $\widehat{g}_{\alpha}$ choose policies from a subset $\hat{\mathcal{G}}_{\alpha}$ defined in Theorem 3.1 and maximize $\widehat{W}_{n}(g)$ as before. Then as a direct consequence of Theorem 3.1, with probability at least $1-\alpha$ this modified rule is guaranteed to not mistakenly choose infeasible eligibility policies.
@@
$\limsup_{n\rightarrow\infty}\sup_{P\in\mathcal{P}}Pr_{P^{n}}\{\hat{\mathcal{G}}_{\alpha}\cap\mathcal{G}^{+}\neq\emptyset\}<\alpha,$ (13)
@@
In practice, $\alpha$ may be set at the conventional level, e.g. 5% for an explicit coverage guarantee. The drawback is that $\widehat{g}_{\alpha}$ can be welfare inefficient. One may therefore let $\alpha_{n}\rightarrow 0$ as sample size increases to achieve both pointwise asymptotically welfare-efficient and pointwise asymptotically feasible for distributions where the constraint is slack at the constrained optimal eligibility policy. Corollary 3.2 formalizes the rate requirement below.
@@
If the data distribution $P$ satisfies Assumption 3.1 and induces a constrained optimal policy at which the constraint is slack, then the modified sample-analog rule
@@
is asymptotically welfare-efficient and asymptotically feasible under $P$ for $\alpha_{n}\rightarrow 0$ at a rate such that $c_{\alpha_{n}}=o(n^{1/2})$.

### v3

@@ 2.3 An impossibility result
Uniformity is a desirable property for a statistical rule, ensuring that its performance guarantees hold uniformly over a class of DGPs, thereby providing robustness to uncertainty about the true DGP. However, in this section, I prove an impossibility result that no statistical rule can be both uniformly asymptotically welfare-efficient and uniformly asymptotically feasible in a sufficiently rich class of distributions $\mathcal{P}$ described in the following Assumptions 2.1-2.3. The intuition is that there exists a sequence $P_{h_{n}}$ that converges to $P_{0}\in\mathcal{P}$, but along the sequence, their corresponding constrained optimum $g^{\ast}_{h_{n}}$ does not converge to the constrained optimum $g^{\ast}_{P_{0}}$ under $P_{0}$. Assumptions 2.2 and 2.3 characterize such point of discontinuity $P_{0}$.
@@
Contiguity. There exists a distribution $P_{0}\in\mathcal{P}$ under which the set $\mathcal{G}_{0}=\left\{g:B(g;P_{0})=k\right\}$ of eligibility policies satisfying the constraint exactly is non-empty. Furthermore, the class of distributions $\mathcal{P}$ includes a sequence of data distributions $\{P_{h_{n}}\}$ contiguous to $P_{0}$, under which for all $g\in\mathcal{G}_{0}$, there exists some $c>0$ such that
@@
Assumption 2.1 assumes there exists a sequence of data distributions $\{P_{h_{n}}\}$ that differ only marginally relative to $P_{0}$. Moreover, the budget function $B(g;P_{h_{n}})$, evaluated at polices that meet the budget constraint exactly under $P_{0}$, approaches $B(g;P_{0})$ from above. In Appendix B.2, I give more primitive assumptions under which Assumption 2.1 is guaranteed to hold, requiring the sequence to be differentiable in quadratic mean at $P_{0}$, and along the sequence, the budget function $B(g;P_{h_{n}})$ is twice continuously differentiable at $P_{0}$ with positive derivatives. These primitive assumptions are relatively weak, and have also been considered in the literature to construct the local parametrization around $P_{0}$, e.g. Hirano and Porter (2009).
@@
Figure 2.1 also highlights the importance of Assumptions 2.2 and 2.3 in driving the impossibility result. Importantly, under $P_{0}$, the budget constraint is binding at the constrained optimal threshold (Assumption 2.2), and increasing the eligibility threshold here has a strictly positive impact on welfare but zero impact on budget (Assumption 2.3). If, instead, increasing the threshold strictly raises both welfare and budget, then Assumption 2.3 is violated. In Section 3.4, I demonstrate that relaxing either Assumption 2.2 or 2.3 gives rise to statistical rules that achieve uniformity within certain subclasses of DGPs. Nonetheless, the impossibility result remains relevant for some real-world policy settings where such assumptions may approximately hold. While the nominal cost of welfare program eligibility may be positive, the Marginal Value of Public Funds (MVPF) framework emphasizes that the relevant cost is the net cost, which incorporates fiscal externalities, such as increased tax revenue if the program increases individuals’ incomes. Specifically, Hendren and Sprung-Keyser (2020) estimate fourteen welfare programs (out of 133) to have negative or zero net cost to the government, which implies these programs “pay for themselves”, aligning with Assumptions 2.2 and 2.3. As a result, the impossibility result suggests that statistical rules designed for such cases may be highly sensitive to small changes in the underlying distributions even in large samples. [^7] As estimated by Hendren and Sprung-Keyser (2020), expanding Medicaid, as in Example 2.1, entails a strictly positive net cost. The impossibility result discussed in this section therefore does not apply to this example. Instead, I use this example to illustrate a new trade-off rule proposed later in Section 4.
@@ 3 Sample-analog rule
I first describe the EWM approach (Kitagawa and Tetenov, 2018) and its direct extension. Since $(\tau,C)$ involves potential outcomes, they are often unobserved and require estimation based on RCT that introduces estimation errors in addition to sampling errors. Section 3.2 describes how to construct individuals’ benefit and cost estimates $(\tau_{i}^{\ast},C_{i}^{\ast})$. To highlight the drawback of the direct extension to EWM, I first consider settings where we observe an experimental data of sample size $n$ where $(\tau,C)$ is directly observable, i.e. $(\tau_{i}^{\ast},C_{i}^{\ast})=(\tau_{i},C_{i})$. The goal of the simplification is to highlight that the non-uniformity I show below can arise from sampling errors alone.
@@
$\widehat{W}_{n}(g)\coloneqq\frac{1}{n}\sum_{i}\tau_{i}^{\ast}\cdot g(X_{i}),\ \widehat{B}_{n}(g)\coloneqq\frac{1}{n}\sum_{i}C_{i}^{\ast}\cdot g(X_{i}).$ (3)
@@ 3.1 Welfare inefficiency of the sample-analog rule
A key insight from Kitagawa and Tetenov (2018) is that without a constraint, the sample-analog rule is uniformly asymptotically welfare-efficient. Unfortunately this intuition does not extend to the current setting where the constraint involves an unknown cost. There are common data distributions under which the amount of welfare loss does not vanish even as the sample size gets larger. Consider a one-dimensional policy class $\mathcal{G}=\{g:g(x)=\mathbf{1}\{x\leq t\}\}$, which is based on thresholds of a one-dimensional continuous characteristic $X$. Suppose the policymakers know benefit is positive for everyone so that welfare function is strictly increasing, and only need to estimate whether a given threshold satisfies a capacity constraint due to imperfect take-up. Furthermore, suppose the experiment sample observes take-up $C_{i}$ so that the only uncertainty arises from sampling errors. Proposition 3.1 shows settings with some zero take-ups satisfy Assumptions 2.2 and 2.3, and the sample-analog rule is neither pointwise asymptotically welfare efficient nor pointwise asymptotically feasible.
@@
One-dimensional threshold and imperfect take-up. Consider a one-dimensional policy class $\mathcal{G}=\{g:g(x)=\mathbf{1}\{x\leq t\}\}$, which includes thresholds for a one-dimensional continuous characteristic $X$. Consider the special case where under distribution $P$, the benefit $\tau>0$ almost surely and the cost $C\in\{0,1\}$ is binary. Suppose $C=0$ for $X\in[\underline{t},\overline{t}]$, but $\operatorname{Pr}_{P}\{C=1\mid X\}\in(0,1)$ otherwise. Suppose further the budget constraint is at $k=\mathbb{E}_{P}[C\cdot\mathbf{1}\{X\leq\underline{t}\}].$ Then $P$ satisfies Assumptions 2.2 and 2.3. As $n\rightarrow\infty$, the sample-analog rule $\widehat{g}_{\text{sample}}$ violates the budget constraint with probability approaching 50%, and there exists an $\epsilon>0$ such that it incurs at least an $\epsilon$ welfare loss with probability approaching 50%:
$\operatorname{Pr}_{P^{n}}\left\{W(g_{P}^{*};P)-W(\widehat{g}_{\text{sample}};P)\geq\epsilon\right\}\rightarrow 50\%.$
@@
The more problematic case is illustrated in Figure 3.1 panel (b), where the sample-analog rule picks a suboptimal threshold because the sample-analog constraint is violated at the constrained optimum $g_{P}^{\ast}$. In the setup of Proposition 3.1, when the sample-analog rule $\widehat{g}_{\text{sample}}$ misses $g_{P}^{\ast}$, it is guaranteed to select a suboptimal policy and therefore it is welfare-inefficient under $P$ even asymptotically. This result relies on the assumption that the welfare function strictly increases when the budget function remains constant in the neighborhood of the budget constraint. This aligns with Assumption 2.3, which, as discussed in Section 2.3, reflect real-world scenarios where some welfare programs have zero or negative net cost to the government. I demonstrate that the same issue extends to broader contexts by employing a simulation calibrated to a real-world DGP from the OHIE in Section 5. If instead the budget function is strictly increasing and violates Assumption 2.3, I show the sample-analog rule is asymptotically welfare-efficient in Proposition C.1 of Appendix C.
@@ 3.2 Estimates for benefit and cost
In the remainder of the paper, I present constructive results that modify the sample-analog rule and propose a new rule. To lay the groundwork, this section outlines the benefit and cost estimates. The appropriate expressions for these estimates depend on the type of observed data. Below I state the estimates formed based an RCT that randomly assigns the eligibility, which is the leading case of Kitagawa and Tetenov (2018). The observed data $\{A_{i}^{\ast}\}_{i=1}^{n}$ consists of i.i.d. observations $A_{i}^{\ast}=(Y_{i},Z_{i},D_{i},X_{i})\in\mathcal{A}^{\ast}$. The distribution of $A_{i}^{\ast}$ is induced by the distribution of $(Y_{1},Y_{0},C,X)$ as in the population, as well as the sampling design of the RCT. Here $D_{i}$ is an indicator for being in the eligibility arm of the RCT, $Y_{i}$ is the observed outcome and $Z_{i}$ is the observed cost of providing eligibility to an individual participating in the RCT. The observed cost is mechanically zero if an individual is not randomized into the eligibility arm. The estimates for $(\tau,C)$ are
$\tau_{i}^{\ast}=\alpha(X_{i},D_{i})\cdot Y_{i},\ C_{i}^{\ast}=\frac{D_{i}}{p(X_{i})}\cdot Z_{i},$ (5)
@@
$\left\{\sqrt{n}\cdot\left(\frac{1}{n}\sum_{i}\tau_{i}^{\ast}\cdot g(X_{i})-W(g;P)\right)\right\}_{g\in\mathcal{G}}$ $\rightarrow_{d}G_{P}^{W}$
$\left\{\sqrt{n}\cdot\left(\frac{1}{n}\sum_{i}C_{i}^{\ast}\cdot g(X_{i})-B(g;P)\right)\right\}_{g\in\mathcal{G}}$ $\rightarrow_{d}G_{P}^{B}$
@@
Appendix B.3 gives primitive assumptions under which Assumption 3.1 is guaranteed to hold for $\widehat{W}_{n}(\cdot)$ and $\widehat{B}_{n}(\cdot)$ constructed using an RCT such as in (5) or an observational study, assuming unconfoundedness and strong overlap. As standard in the literature, I need to restrict the complexity of the policy class $\mathcal{G}$. The policy class of income thresholds considered in this paper is in fact a rather simple class with VC-dimension $d+1$ where $d$ is the number of different thresholds as in (2).
@@ 3.3 Modification that ensures feasibility
Let $\hat{\mathcal{G}}=\{g\in\mathcal{G}:\widehat{B}_{n}(g)\leq k\}$ denote the set of policies that $\widehat{g}_{\text{sample}}$ can choose from, which contains policies that do not violate the sample-analog of the budget constraint. Unsurprisingly, for any finite sample, $\hat{\mathcal{G}}$ can always contain infeasible policies, sometimes of sizable budget violations. The probability that $\hat{\mathcal{G}}$ contains a policy that violates the population budget constraint by a fixed amount $c$ is as follows:
$\operatorname{Pr}_{P^{n}}\left\{\exists g:g\in\hat{\mathcal{G}}\text{ and }B(g;P)>k+c\right\}$ (6)
@@
A simple modification to the sample-analog rule can reduce the probability of selecting infeasible policies. Instead of $\hat{\mathcal{G}}$, let $\widehat{g}_{\alpha}$ choose policies from a subset $\hat{\mathcal{G}}_{\alpha}$ defined in Theorem 3.1 and maximize $\widehat{W}_{n}(g)$ as before. Then as a direct consequence of Theorem 3.1, with probability at least $1-\alpha$ this modified rule is guaranteed to not mistakenly choose infeasible eligibility policies. In practice, $\alpha$ may be set at the conventional level, e.g. 5%. Then effectively this modification first forms a uniform upper confidence band for the costs of all the policies with asymptotic coverage at least 95% and $\hat{\mathcal{G}}_{\alpha}$ collects only the policies for which the upper bound on cost is below the threshold.
@@
$\limsup_{n\rightarrow\infty}\sup_{P\in\mathcal{P}}\operatorname{Pr}_{P^{n}}\{\hat{\mathcal{G}}_{\alpha}\cap\mathcal{G}^{+}\neq\emptyset\}<\alpha,$
@@
3.4 Modification that ensures uniformity in subclasses of DGPs
Although $\widehat{g}_{\alpha}$ achieves asymptotic feasibility uniformly over a wide class of DGPs, it may lead to welfare inefficiencies under certain DGPs. Below I show that if one lets $\alpha_{n}\rightarrow 0$ as sample size increases, at a rate such that $c_{\alpha_{n}}=o(n^{1/2})$, then the modified sample-analog rule $\widehat{g}_{\alpha_{n}}$ is both asymptotically welfare-efficient and asymptotically feasible uniformly over two subclasses of distributions.
@@ 3.4 Modification that ensures uniformity in subclasses of DGPs
Consider a family of distributions $\mathcal{P}$ such that for all $P\in\mathcal{P}$, Assumption 3.1 holds and the budget constraint is uniformly slack under the constrained optimal policy. That is, there exists $\delta>0$ such that
$B(g_{P}^{\ast};P)\leq k-\delta\quad\text{for all }P\in\mathcal{P},$
so that Assumption 2.2 is violated for every $P\in\mathcal{P}$. Then the modified sample-analog rule
@@
is asymptotically welfare-efficient and asymptotically feasible uniformly under $\mathcal{P}$ for $\alpha_{n}\rightarrow 0$ at a rate such that $c_{\alpha_{n}}=o(n^{1/2})$.
Corollary 3.3.
Let the policy class $\mathcal{G}$ be based on continuously distributed covariates $X$ and parameterized by a finite-dimensional threshold parameter $\theta$. Consider a family of distributions $\mathcal{P}$ such that for all $P\in\mathcal{P}$, Assumption 3.1 holds. Moreover, for every $P\in\mathcal{P}$, both the conditional benefit $\mathbb{E}_{P}[\tau\mid X]$ and the conditional cost $\mathbb{E}_{P}[C\mid X]$ are strictly positive. In particular, there exists $\underline{c}>0$ such that
$\mathbb{E}_{P}[C\mid X]\geq\underline{c}\quad\text{for all }P\in\mathcal{P},$
so that Assumption 2.3 is violated for every $P\in\mathcal{P}$. Then the modified sample-analog rule $\widehat{g}_{\alpha_{n}}$, as defined in (8), is asymptotically welfare-efficient and asymptotically feasible uniformly under $\mathcal{P}$.
Note that the above results do not contradict the impossibility result in Section 2.3, which shows lack of uniformity over a broader class of distributions, including those satisfying Assumptions 2.2 and 2.3. Specifically, Corollary 3.2 establishes uniformity within a subclass of distributions that violate Assumption 2.2, while Corollary 3.3 establishes uniformity within a subclass of distributions that violate Assumption 2.3.